OpenStudy (anonymous):

What is the quotient: (6x4 – 15x3 – 2x2 – 10x – 4) ÷ (3x2 + 2) ?

OpenStudy (angela210793):

24-45-4+40 20 ------------=------=5/2 8 8

OpenStudy (anonymous):

\[\frac{6 x^4-15 x^3-2 x^2-10 x-4}{3 x^2+2}=2 x^2-5 x-2 \]\[\left(3 x^2+2\right)\left(2 x^2-5 x-2\right)=6 x^4-15 x^3-2 x^2-10 x-4 \]

OpenStudy (angela210793):

those were exponents?

OpenStudy (anonymous):

Most of the math expressions put forth on this site are Ambiguous and this is one of them. Pending feedback, I don't know if my interpretation is correct or not.

OpenStudy (angela210793):

:)